OpenStudy (anonymous):

what is the y intercept of 2x-4=3y

jimthompson5910 (jim_thompson5910):

to find the y-intercept, you plug in x = 0 and solve for y

jimthompson5910 (jim_thompson5910):

2x-4=3y 2(0)-4=3y 0 - 4 = 3y -4 = 3y ... ... ... y = ???

OpenStudy (anonymous):

profit